OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = –3x + 9 A. slope: 9, y-intercept: –3 B. slope: 1, y-intercept: –3 C. slope: 3, y-intercept: –9 D. slope: –3, y-intercept: 9

OpenStudy (anonymous):

d

OpenStudy (anonymous):

slope=-3, y-int=9

OpenStudy (anonymous):

y=mx+b m=slope b=y-int

OpenStudy (anonymous):

cna you help me on one more?

OpenStudy (anonymous):

What is the slope-intercept form of this equation? 3x + 9y = 18

OpenStudy (anonymous):

Slope-1/3 y-intercept 2

OpenStudy (ipwnbunnies):

Don't just tell her the answer. Explain it.

OpenStudy (anonymous):